# Another look at inference after prediction

### v2

@@ Abstract
Prediction-based (PB) inference is increasingly used in applications where the outcome of interest is difficult to obtain, but its predictors are readily available. Unlike traditional inference, PB inference performs statistical inference using a partially observed outcome and a set of covariates by leveraging a prediction of the outcome generated from a machine learning (ML) model. Motwani and Witten (2023) recently revisited two innovative PB inference approaches for ordinary least squares. They found that the method proposed by Wang et al. (2020) yields a consistent estimator for the association of interest when the ML model perfectly captures the underlying regression function. Conversely, the prediction-powered inference (PPI) method proposed by Angelopoulos et al. (2023a) yields valid inference regardless of the model’s accuracy. In this paper, we study the statistical efficiency of the PPI estimator. Our analysis reveals that a more efficient estimator, proposed 25 years ago by Chen and Chen (2000), can be obtained by simply adding a weight to the PPI estimator. We also contextualize PB inference with methods from the economics and statistics literature dating back to the 1960s. Our extensive theoretical and numerical analyses indicate that the Chen and Chen (CC) estimator offers a balance between robustness to ML model specification and statistical efficiency, making it the preferred choice for use in practice.
Keywords: Missing data, prediction-based inference, semi-parametric inference
@@ 1 Introduction
In the rapidly evolving landscape of data science, accurately predicting an outcome is often just an initial step in scientific inquiry. Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to directly measure, and then drawing statistical inferences from the ML model’s predictions (Hoffman et al., 2024; Fan et al., 2024; Gao et al., 2024; Miao and Lu, 2024). This prediction-based (PB) inference scenario was introduced by Wang et al in the context of global health and genomics applications, and is comprised of two steps outlined in Figure 1 (Wang et al., 2020). First, the outcome of interest is imputed using a prediction from a pre-trained model. Then, the predictions are used to draw statistical inference on the parameter of interest, such as the association between an outcome and a set of covariates.
@@
Figure 1: Prediction-based (PB) inference. $Y$ is the outcome of interest, which is only partially observed, ${\bf{Z}}$ is a set of predictors of $Y$, $\widehat{Y}$ is the prediction of $Y$ based on ${\bf{Z}}$, and ${\bf{X}}$ is a set of covariates. The focus of scientific inquiry is on the association between ${\bf{X}}$ and $Y$. Traditional statistical inference relies solely on the data on ${\bf{X}}$ and $Y$ while PB inference incorporates $\widehat{Y}$ to maximize the use of all available data. Dark red represents labeled observations while light red represents unlabeled observations.
@@
In the last several years, applications of PB inference have arisen in diverse fields, spanning genetics, biology, and medicine. For example, population biobanks, such as the UK Biobank ($N\approx$ 500K individuals), provide an unprecedented data source for high-throughput genetic research. However, many traits and disease outcomes are only partially observed due to the time and cost of measurement. ML models are commonly used to predict missing outcomes based on fully observed baseline data and the predictions are utilized for genetic discovery (Tong et al., 2020; Alipanahi et al., 2021; Cosentino et al., 2023; Dahl et al., 2023; An et al., 2023; McCaw et al., 2024; Liu et al., 2024; Miao et al., 2024). Similarly, in the field of structural biology, AlphaFold provides highly accurate structural predictions without the need for expensive and labor-intensive experiments for proteomic studies (Jumper et al., 2021). In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review by experts, enabling downstream epidemiological and clinical studies (Yang et al., 2022; Zhou et al., 2022; Alsentzer et al., 2023).
@@
Given the relevance of PB inference, Motwani and Witten (2023) recently revisited two innovative methods in this area. In the context of least squares estimation, they showed that the approach proposed by Wang et al. (2020) provides a consistent estimator for the association between an outcome and a set of covariates only in restrictive scenarios. For example, when the pre-trained model perfectly captures the regression function. This is a significant practical limitation as it is essential that PB inference methods yield valid inference in the likely scenario that the prediction model is imperfect. In contrast, the prediction powered inference (PPI) approach proposed by Angelopoulos et al. (2023a) possesses this necessary property and yields valid inference regardless of the quality of the model.
@@
Here we build on the work on Motwani and Witten to study the statistical efficiency of the PPI method. We show that a more efficient estimator can be obtained by simply adding a weight to the PPI estimator, an approach that was proposed by Chen & Chen (CC) more than 25 years ago in the context of double sampling designs (Chen and Chen, 2000). We begin by formally introducing the PB inference problem and reviewing the contributions of Motwani and Witten in Section 2. We then study the asymptotic properties of the PPI estimator and introduce the CC estimator in Section 3. Our key finding is that the CC estimator is optimal with respect to statistical efficiency in a class of estimators to which the PPI estimator belongs. We also point to earlier literature from the statistics and economics literature dating back to the 1960s that provides important historical context for the PB inference problem in Section 4. In Section 5, we evaluate our findings with extensive simulation studies. Our results show that the CC estimator offers a balance between robustness to ML model specification and statistical efficiency, making it the preferred choice for use in practice. We close with additional discussion in Section 6.
@@ 2.1 Notation and Assumptions
We focus on the case in which $\mathscr{L}$ and $\mathscr{U}$ are independent and identically distributed samples from the population of interest. More complex scenarios have been the focus of recent work (Angelopoulos et al., 2023a; Zrnic and Candès, 2024a). Additionally, $\hat{f}$ is assumed to be pre-trained with a dataset that is independent of $\mathscr{L}$ and $\mathscr{U}$ and is therefore treated as “fixed” (i.e., the variation in estimating $\hat{f}$ is not of interest in downstream inference). In fact, $\hat{f}$ may be a black box function in that its training data and specification are not be available to the users. Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been proposed in Zrnic and Candès (2024b).
@@ 2.2 Parameter of interest
Following Motwani and Witten (2023), we consider least squares estimation and aim to make inference on the population parameter
@@ 2.3 Motivation for PB inference
The need for PB inference is highlighted by limitations of “naive” inference, in which inference on $\widehat{Y}$ is performed in lieu of inference on $Y$ (Wang et al., 2020). The parameter targeted by naive inference is
@@ 2.4 PB inference methods
Motwani and Witten revisited two innovative strategies for PB inference. The first is the proposal of Wang et al. (2020), which is a two-step procedure to incorporate $\widehat{Y}$ into estimation. First, $\mathscr{L}$ is used to model the relationship between $Y$ and $\widehat{Y}$ to address potential inaccuracies in the predictions. Then, this model is used together with $\mathscr{U}$ to make inference on $\bm{\beta}^{*}$, with analytical or bootstrap approaches described. In their illuminating work, Motwani and Witten show that the estimator of Wang et al is consistent for $\bm{\beta}^{*}$ under the strong and unrealistic assumption that the ML model captures the true regression function. In other scenarios, the Wang et al estimator may not be consistent for $\bm{\beta}^{*}$ even while its standard error converges to 0 as $(n_{\text{\tiny{lab}}},n_{\text{\tiny{unlab}}})\to\infty$. Consequently, the familiar Wald statistic will not converge in distribution under the null, resulting in type I error that is not controlled at the nominal level and confidence intervals with inadequate coverage.
Motwani and Witten also studied the PPI method, which has the crucial property that inference remains valid regardless of the quality of the predictions. Stated otherwise, the PPI method does not require that $\hat{f}$ capture the true regression function for the estimator to maintain consistency or to yield reliable hypothesis testing and interval estimation. The proposed estimator utilizes a common statistical technique involving the augmentation of an initial consistent estimator for the parameter of interest with an additional term that converges in probability to 0. This approach dates back to at least Robins et al. (1994), and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem (e.g., Chen and Chen (2000); Tsiatis (2006); Tian et al. (2012); Wang and Wang (2015)). We return to this point in Section 4.
@@
where $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$ is the solution to $\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\boldsymbol{0}$ [^1] Angelopoulos et al. (2023a) express $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ where $\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ is referred to as a “rectifier.” The rectifier serves the purpose of correcting potential bias in $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ due to inaccuracies in $\widehat{Y}$ for $Y$. We reframe the PPI method to study its statistical efficiency.. In contrast to the approach of Wang et al, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is consistent for $\bm{\beta}^{*}$ and $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*})$ converges weakly to a zero-mean multivariate normal distribution regardless of the quality of the prediction. The proposed method therefore asymptotically controls the type I error for Wald-based hypothesis testing and provides confidence intervals with the nominal coverage level.
@@ 3 Efficiency analysis
where $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ is the solution to $\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\boldsymbol{0}$. The only difference between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that all of the data, rather than the unlabeled data, is used to estimate $\bm{\gamma}^{*}$ in the augmentation term (i.e., $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ is replaced by $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$). When $\pi\approx 0$, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}\approx\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ so that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}\approx\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$. In general, it can be shown that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ gains over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when $\widehat{Y}$ is informative of $Y$ (see Appendix B, specifically Section B.3, for a detailed analysis). Additionally, estimators of this form are commonly used in the semi-parametric inference literature, which we draw upon in the subsequent section in our efficiency analysis.
@@ 3.1 Asymptotic analysis of the PPI estimator
Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$)
@@
Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
Let $\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}=\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$, $\boldsymbol{\Sigma}_{Y,\widehat{Y}}=\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$, $\epsilon_{Y}=Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$ and $\epsilon_{\widehat{Y}}=\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}$. The difference in the asymptotic variance-covariance (aVar) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=n\left[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})-\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})\right]=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$. A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
@@
The sufficient condition in Corollary 1.1 coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on ${\bf{X}}$. This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al. (2023a) and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix A.4). Here we consider a simple example to demonstrate when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ under homoskedasticity.
Example 1 (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ under homoskedasticity)
@@
where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the correlation of the residuals for $Y$ and $\widehat{Y}$ given ${\bf{X}}$.
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ therefore gains efficiency if and only if the correlation of the residuals is at least half of the ratio of residual standard deviations of $\widehat{Y}$ to that of $Y$. Put another way, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ improves over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with respect to efficiency when $\widehat{Y}$ is a strong predictor for $Y$ after adjusting for ${\bf{X}}$, though $\widehat{Y}$ does not necessarily need to be derived from a ML model that captures the true regression function. The improvement is stronger with increasing correlation of the residuals between $Y$ and $\widehat{Y}$ and decreasing variability in the residuals of $\widehat{Y}$.

### v3

@@ Abstract
From structural biology to epidemiology, predictions from machine learning (ML) models increasingly complement costly gold-standard data to enable faster, more affordable, and scalable scientific inquiry. In response, prediction-based (PB) inference has emerged to accommodate statistical analysis using a large volume of predictions together with a small amount of gold-standard data. The goals of PB inference are two-fold: (i) to mitigate bias from errors in predictions and (ii) to improve efficiency relative to traditional inference using only the gold-standard data. Motwani and Witten [2023] recently revisited two key PB inference approaches and found that only one method, Prediction-powered Inference (PPI) proposed by Angelopoulos et al. [2023a], achieves (i). In this paper, we find that PPI does not achieve (ii). We revisit the double sampling literature and show that, with a simple modification, PPI can be adjusted to provide theoretically justified improvements in efficiency. We also contextualize PB inference with economics and statistics literature dating back to the 1960s to highlight the utility of classical methods in this contemporary problem. Our extensive theoretical analyses, along with an analysis of UK Biobank data, indicate that our proposal effectively mitigates bias and improves efficiency, making it preferable for use in practice.
Keywords: Machine learning, measurement error, missing data, prediction-based inference, semi-parametric inference
@@ 1 Introduction
In the rapidly evolving landscape of data science, accurately predicting an outcome is often just an initial step in scientific inquiry. Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to directly measure, and then drawing statistical inferences from the ML model’s predictions [Hoffman et al., 2024, Fan et al., 2024, Gao et al., 2024, Miao and Lu, 2024]. This prediction-based (PB) inference scenario was introduced by Wang et al in the context of global health and genomics applications, and is comprised of two steps outlined in Figure 1 [Wang et al., 2020]. First, the outcome of interest is imputed using a prediction from a pre-trained model. Then, the predictions are used to draw statistical inference on the parameter of interest, such as the association between an outcome and a set of covariates, together with a small amount of gold-standard data  [^1] PB inference has focused on outcomes that are difficult to measure. Recent work has considered predicted covariates (e.g., Kluger et al. [2025]).. In contrast to traditional inference that relies solely on gold-standard data, PB inference aims to leverage the predictions to improve statistical efficiency without introducing bias.
@@
Figure 1: Prediction-based (PB) inference. $Y$ is the outcome of interest, which is only partially observed, ${\bf{Z}}$ is a set of predictors of $Y$, $\widehat{Y}$ is the prediction of $Y$ based on ${\bf{Z}}$, and ${\bf{X}}$ is a set of covariates. The focus of scientific inquiry is on the association between $Y$ and ${\bf{X}}$. Traditional statistical inference relies solely on the data on $Y$ and ${\bf{X}}$ while PB inference incorporates $\widehat{Y}$ to maximize the use of all available data. Dark red represents labeled observations while light red represents unlabeled observations.
@@
In the last several years, applications of PB inference have arisen in diverse fields, spanning genetics, biology, and medicine. For example, population biobanks, such as the UK Biobank ($N\approx$ 500K individuals), provide an unprecedented data source for high-throughput genetic research. However, many traits and disease outcomes are only partially observed due to the time and cost of measurement. ML models are commonly used to predict missing outcomes based on fully observed baseline data and the predictions are utilized for genetic discovery [Tong et al., 2020, Alipanahi et al., 2021, Cosentino et al., 2023, Dahl et al., 2023, An et al., 2023, McCaw et al., 2024, Liu et al., 2024, Miao et al., 2024]. Similarly, in the field of structural biology, AlphaFold provides structural predictions without the need for expensive and labor-intensive experiments for proteomic studies [Jumper et al., 2021, Gan et al., 2022]. In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review by experts, enabling downstream epidemiological and clinical studies [Yang et al., 2022, Zhou et al., 2022, Alsentzer et al., 2023].
@@
Given the relevance of PB inference, Motwani and Witten [2023] recently revisited two innovative methods in this field. In the context of least squares estimation, they showed that the approach proposed by Wang et al. [2020] provides a consistent estimator for the association between an outcome and a set of covariates only in restrictive scenarios. For example, when the pre-trained model perfectly captures the regression function. This is a significant practical limitation as it is essential that PB inference methods yield valid inference in the likely scenario that the prediction model is imperfect. In contrast, the prediction powered inference (PPI) approach proposed by Angelopoulos et al. [2023a] possesses this necessary property and yields valid inference regardless of the quality of the model. This finding is critical given that the PPI paper has been cited over 100 times in its first year of publication in Science.
@@
In this paper, we study the statistical efficiency of the PPI estimator and show that it can be outperformed by traditional inference. We propose a more efficient estimator that can be obtained by simply adding a weight to the PPI estimator, an approach introduced by Chen & Chen (CC) 25 years ago in the context of double sampling designs [Chen and Chen, 2000]. We begin by formally introducing the PB inference problem and reviewing the contributions of Motwani and Witten in Section 2. We then study the asymptotic properties of the PPI estimator and introduce the CC estimator in Section 3. Our key finding is that the CC estimator is optimal with respect to statistical efficiency in a class of estimators to which the PPI estimator belongs. We also review modern and classical literature from statistics and economics dating back to the 1960s that provides important context for the PB inference problem in Section 4. In Section 5, we evaluate our findings with extensive simulation studies. In Section 6, we study the association between android fat mass and clinical and demographic covariates in the South Asian population in the UK Biobank to demonstrate the practical performance of our method. We close with additional discussion in Section 7. All in all, our theoretical and numerical results show that the CC estimator offers a balance between robustness to ML model specification and statistical efficiency, making it the preferred choice for use in PB inference problems.
@@ 2.1 Notation and Assumptions
We focus on the case in which $\mathscr{L}$ and $\mathscr{U}$ are independent and identically distributed samples from the population of interest. More complex scenarios have been the focus of recent work [Angelopoulos et al., 2023a, Zrnic and Candès, 2024a]. Additionally, $\hat{f}$ is assumed to be pre-trained with a dataset that is independent of $\mathscr{L}$ and $\mathscr{U}$ and is therefore treated as “fixed” (i.e., the variation in estimating $\hat{f}$ is not of interest in downstream inference). In fact, $\hat{f}$ may be a black box function in that its training data and specification are not be available to the users. Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been proposed by Zrnic and Candès [2024b].
@@ 2.2 Parameter of interest
Following Motwani and Witten [2023], we consider least squares estimation and aim to make inference on the population parameter
@@ 2.3 Motivation for PB inference
The need for PB inference is highlighted by limitations of “naive” inference, in which inference on $\widehat{Y}$ is performed in lieu of inference on $Y$ [Wang et al., 2020]. The parameter targeted by naive inference is
@@ 2.4 PB inference methods
Motwani and Witten revisited two innovative strategies for PB inference. The first is the proposal of Wang et al. [2020], which is a two-step procedure to incorporate $\widehat{Y}$ into estimation. First, $\mathscr{L}$ is used to model the relationship between $Y$ and $\widehat{Y}$ to address potential inaccuracies in the predictions. Then, this model is used together with $\mathscr{U}$ to make inference on $\bm{\beta}^{*}$ with either analytical or bootstrap approaches. In their illuminating work, Motwani and Witten show that the estimator of Wang et al is consistent for $\bm{\beta}^{*}$ under the strong and unrealistic assumption that the ML model captures the true regression function. In other scenarios, the Wang et al estimator may not be consistent for $\bm{\beta}^{*}$ even while its standard error converges to 0 as $(n_{\text{\tiny{lab}}},n_{\text{\tiny{unlab}}})\to\infty$. Consequently, the familiar Wald statistic will not converge in distribution under the null, resulting in type I error that is not controlled at the nominal level and confidence intervals with inadequate coverage.
Motwani and Witten also studied the PPI method, which has the crucial property that inference remains valid regardless of the quality of the predictions. Stated otherwise, the PPI method does not require that $\hat{f}$ capture the true regression function for the estimator to maintain consistency or to yield reliable hypothesis testing and interval estimation. The proposed estimator utilizes a common statistical technique involving the augmentation of an initial consistent estimator for the parameter of interest with an additional term that converges in probability to 0. This approach dates back to at least Robins et al. [1994], and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem (e.g., Chen and Chen [2000], Tsiatis [2006], Tian et al. [2012], Wang and Wang [2015]). We return to this point in Section 4.
@@
where $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$ is the solution to $\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$ [^2] Angelopoulos et al. [2023a] express $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ where $\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ is referred to as a “rectifier.” The rectifier serves the purpose of correcting potential bias in $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ due to inaccuracies in $\widehat{Y}$ for $Y$. We reframe the PPI method to study its statistical efficiency.. In contrast to the approach of Wang et al, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is consistent for $\bm{\beta}^{*}$ and $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*})$ converges weakly to a zero-mean multivariate normal distribution regardless of the quality of the prediction. The proposed method therefore asymptotically controls the type I error for Wald-based hypothesis testing and provides confidence intervals with the nominal coverage level.
@@ 3 Efficiency analysis
where $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ is the solution to $\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$. The only difference between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that all of the data, rather than the unlabeled data, is used to estimate $\bm{\gamma}^{*}$ in the augmentation term (i.e., $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ is replaced by $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$). When $\pi\approx 0$, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}\approx\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ so that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}\approx\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$. In general, it can be shown that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ gains over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when $\widehat{Y}$ is informative of $Y$ (see Appendix B, specifically Section B.3, for a detailed analysis). Additionally, estimators of this form are commonly used in the semi-parametric inference literature, which we draw upon in the subsequent section in our efficiency analysis.
@@ 3.1 Asymptotic analysis of the PPI estimator
Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$).
@@
Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
Let $\epsilon_{Y}=Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$, $\epsilon_{\widehat{Y}}=\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}=\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$, and $\bm{\Sigma}_{Y,\widehat{Y}}=\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$. The difference in the asymptotic variance-covariance (aVar) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=n\left[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})-\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})\right]=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$. A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
@@
The sufficient condition in Corollary 1.1 coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on ${\bf{X}}$. This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al. [2023a] and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix A.4). Here we consider a simple example to demonstrate when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ under homoskedasticity.
Example 1 (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ under homoskedasticity).
@@
where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the correlation of the residuals for $Y$ and $\widehat{Y}$ given ${\bf{X}}$. Put another way, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ improves over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with respect to efficiency when $\widehat{Y}$ is a strong predictor for $Y$ after adjusting for ${\bf{X}}$, though $\widehat{Y}$ does not necessarily need to be derived from a ML model that captures the true regression function. The improvement is stronger with increasing correlation of the residuals between $Y$ and $\widehat{Y}$ and decreasing variability in the residuals of $\widehat{Y}$.